OpenStudy (goformit100):

The length of the bridge, which a train 130 metres long and travelling at 45 km/hr can cross in 30 seconds, is: A. 200 m B. 225 m C. 245 m D. 250 m

OpenStudy (goformit100):

@Noemi95

OpenStudy (anonymous):

Let x be the length of the bridge. The engine would first need to cross the bridge, and then another 130 m (length of the train) so that the entire train crosses the bridge. So distance to be covered = x + 130 And all this needs to be done in 30 seconds Speed = (x + 130) / 30 Given speed = 45 km/hr = 12.5 m/s (x + 130) / 30 = 12.5 x + 130 = 375 x = 245 m